OpenStudy (anonymous):

\[simplify 5/2 \sqrt{48}\]

OpenStudy (ash2326):

\[\frac 52 \sqrt {48}\] Is this your question?

OpenStudy (anonymous):

yes

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \sqrt{48} = \sqrt{2 * 2 * 2 * 2 * 3} = 4\sqrt{3} }\) Continue?

OpenStudy (ash2326):

@odalys79 Now use the hint provided by @ParthKohli to simplify the expression

Parth (parthkohli):

Also, \(\Large \color{Black}{\Rightarrow z \times x\sqrt{y} = zx\sqrt{y} }\) No more hints xD

OpenStudy (anonymous):

so now i have to figure out the fraction?

OpenStudy (anonymous):

\[10 4\sqrt{3}\]

OpenStudy (anonymous):

\[10 4\sqrt{3} \]

OpenStudy (anonymous):

10x12=120